You rev your car's engine and watch the tachometer climb steadily from 1400 rpm to 5400 rpm in 3.1 s. (a) What is the angular acceleration of the engine? rad/s2 (b) What is the tangential acceleration of a point on the edge of the engine's 3.5 cm diameter crankshaft? m/s2 (c) How many revolutions does the engine make during this time? rev
Answer #1

a)

Angular acceleration is

\alpha =\frac{\Delta w}{\Delta t}

= (5400-1400)(2pi/60) / 3.1 = 135.12 rad/s^2

b)

tangential acceleration is

a_{t}=\alpha r

= 135.12 * 3.5*10^(-2) / 2 = 2.365 m/s^2

c)

\Delta \Theta =w_{0}t+\frac{\alpha t^{2}}{2}

where w0 is initial speed = 1400*2pi/60 = 146.61 m/s

\Delta \Theta = 146.61*3.1 + 135.12*(3.1^2)/2

= 1103.74 rad = 1103.74/2pi = 175.67 rev